In concentration of Aqueous solutions find mass fraction and Mole Fraction using the volume fraction of 0.8629, Specific Gravity of solute of 0.4040 and molecular weight of solute of 59.00



1
Expert's answer
2021-10-04T05:10:58-0400

Volume fraction "= \frac{Mass\; fraction}{Specific\; gravity}"

Mass fraction "= Volume \;fraction \times Specific\; gravity"

Mass fraction "= 0.8629 \times 0.4040 = 0.3487"

Mole fraction "= \frac{Mass \;fraction}{Molecular\; weight}"

Mole fraction "= \frac{0.6487}{59.00}= 0.0059"